# Computationally Efficient Electromagnetic Transient Power System Studies using Bayesian Optimization

### v2

@@ Abstract
Keywords:
Bayesian Optimization, Classification, Electromagnetic Transients
@@ II Automation Settings
Fig. 1: Workflow for Bayesian Optimization (a), Active Learning (b), Active Search (c) and Active Discovery (d).
@@ V Case Study
In this case study, the automation framework is applied to the assessment of transformer stresses, i.e., overvoltages, during energization in the vicinity of underground cables. The transformer energization problem is extensively discussed in [2]. The application study borrows the parameters from [20] in which the case study is outlined.
@@
The case study makes use of a transformer connected to a reduced network as shown in [20]. The transformer is modeled using the classical modeling approach with saturation modeled across the magnetizing inducance. The grid is modeled by a simple RLC-network, as also proposed by a system operator to mimic a transformer energized through a long cable [22]. The transformer parameters for the case study have been taken from [20]. The RLC parameters are selected as $R=1.32\leavevmode\nobreak\ \Omega$, $L=50$ mH, $C=50.6$ $\mu$F, corresponding to a resonance peak in the input impedance of amplitude 750 $\Omega$ at 100 Hz. As output of the simulation, the phase-to-phase voltage between phases a and b is chosen. The simulation model is implemented in EMT-type software, and executed with a time-step of 50 $\mu$s for a total simulation duration of 2.5 s.
@@
Fig. 2: Electrical circuit for transformer energization analysis.
V-A Brute Force Benchmark Solution
@@
Fig. 3: Brute Force Benchmark Solution
V-B Maximum Overvoltage (One Dimensional Input)
@@
Fig. 4: Optimal value against function evaluations (1D Input)
Fig. 5: Surrogate model and benchmark for BO
V-C Maximum Overvoltage (Three-Dimensional Input)
@@
The Bayesian Optimizer is able to find the global optimum for the three-dimensional problem with a total of 70 function evaluations (Fig. 6). The optimization is terminated once the stopping criterion, i.e., the fixed budget of function evaluations, is reached. Compared with the brute force benchmark solution, the BO reduces the number of EMT simulations by a factor of 100. The ability of the optimizer to find the global optimum depends on the number of samples used to initialize the model. In this case, the optimizer has been initialized with 50 samples, indicated with a cross, while using 20 iterations for the actual optimization, indicated with a dot. If initialized with an insufficient number of samples, the optimizer may output a local optimum. It should be noted that the optimum differs from the previous case due to the additional aleatoric input variables.
Fig. 6: Optimal value against function evaluations (3D Input)
@@
Fig. 7: Maximum overvoltage for different optimization methods, showing the output for 10 randomly initiated attempts with a simulation budget of 100 function evalutions
@@
Fig. 8: Optimality gap against number of function evaluations for different optimization methods for best and worst case of each of the 10 optimization attempts
V-D Binary Classification
@@
Fig. 9: Surrogate model and benchmark for binary classification task

### v3

@@ Abstract
Index Terms:
Bayeisan Optimization, Electromagnetic Transients, Classification
@@ II Automation Settings
Figure 1: Workflow for Bayesian Optimization (a), Active Learning (b), Active Search (c) and Active Discovery (d).
@@ V Case Study
The automation framework is applied to the assessment of transformer stresses during energization in the vicinity of underground cables. The application study borrows the parameters from [20] in which the case study is outlined.
V-A Transformer Energization
This case studies the overvoltages experienced during energization when resonances are matching harmonic frequencies, e.g., when energizing transformers in the vicinity of high-voltage cables. The transformer energization problem is extensively discussed in [2].
@@ V-A Transformer Energization
The case study makes use of a transformer connected to a reduced network as shown in [20]. The transformer is modeled using the classical modeling approach with saturation modeled across the magnetizing inducance. The grid is modeled by a simple RLC-network, as also proposed by a system operator to mimic a transformer energized through a long cable [22]. The transformer parameters for the case study have been taken from [20]. The RLC parameters are selected as $R=1.32~{}\Omega$, $L=50$ mH, $C=50.6$ $\mu$F, corresponding to a resonance peak in the input impedance of amplitude 750 $\Omega$ at 100 Hz. As output of the simulation, the phase-to-phase voltage between phases a and b is chosen. The simulation model is implemented in EMT-type software, and executed with a time-step of 50 $\mu$s for a total simulation duration of 2.5 s.
@@
Figure 2: Electrical circuit for transformer energization analysis.
V-A1 Brute Force Benchmark Solution
@@
Figure 3: Brute Force Benchmark Solution
V-A2 Maximum Overvoltage (One Dimensional Input)
@@
Figure 4: Optimal value against function evaluations (1D Input)
Figure 5: Surrogate model and benchmark for BO
V-A3 Maximum Overvoltage (Three-Dimensional Input)
@@
The Bayesian Optimizer is able to find the global optimum for the three-dimensional problem with a total of 70 function evaluations (Fig. 6). Compared with the brute force benchmark solution, the BO reduces the number of EMT simulations by a factor of 100. The ability of the optimizer to find the global optimum depends on the number of samples used to initialize the model. In this case, the optimizer has been initialized with 50 samples, indicated with a cross, while using 20 iterations for the actual optimization, indicated with a dot. If initialized with an insufficient number of samples, the optimizer may output a local optimum. It should be noted that the optimum differs from the previous case due to the additional aleatoric input variables.
Figure 6: Optimal value against function evaluations (3D Input)
@@
Figure 7: Maximum overvoltage for different optimization methods, showing the output for 10 randomly initiated attempts with a simulation budget of 100 function evalutions
@@
Figure 8: Optimality gap against number of function evaluations for different optimization methods for best and worst case of each of the 10 optimization attempts
V-A4 Binary Classification
@@
Figure 9: Surrogate model and benchmark for binary classification task